Show that C = [T1/2]/[R*ln(2)], where T1/2 is the half-life (time interval required for the voltage/or charge to fall to one-half of its initial value) and RC time constant is τ = R*C.
Answer #1

2^t = e^T1/2

t*ln(2) = T1/2

RC*ln(2) = T1/2

C = T1/2 / [R*ln(2)]